# HyDeFuse: Provably Convergent Denoiser-Driven Hyperspectral Fusion

### v1

@@ Abstract
Hyperspectral (HS) images provide fine spectral resolution but have limited spatial resolution, whereas multispectral (MS) images capture finer spatial details but have fewer bands. HS-MS fusion aims to integrate HS and MS images to generate a single image with improved spatial and spectral resolution. This is commonly formulated as an inverse problem with a linear forward model. However, reconstructing high-quality images using the forward model alone is challenging, necessitating the use of regularization techniques. Over the years, numerous methods have been developed, including wavelets, total variation, low-rank models, and deep neural networks. In this work, we investigate the paradigm of denoiser-driven regularization, where a powerful off-the-shelf denoiser is used for implicit regularization within an iterative algorithm. This has shown much promise but remains relatively underexplored in hyperspectral imaging. Our focus is on a crucial aspect of denoiser-driven algorithm — ensuring convergence of the iterations. It is known that powerful denoisers can produce high-quality reconstructions, but they are also prone to instability and can cause the iterations to diverge. The challenge is to come up with denoisers that come with a convergence guarantee. In this work, we consider a denoiser-driven fusion algorithm, HyDeFuse, which leverages a class of pseudo-linear denoisers for implicit regularization. We demonstrate how the contraction mapping theorem can be applied to establish global linear convergence of HyDeFuse. Additionally, we introduce enhancements to the denoiser that significantly improve the performance of HyDeFuse, making it competitive with state-of-the-art techniques. We validate our theoretical results and present fusion results on publicly available datasets to demonstrate the performance of HyDeFuse.
@@ 1 Introduction
Hyperspectral cameras capture images of a scene across a continuous range of spectral bands. The spectral information is useful in applications such as navigation, surveillance, object identification, and medical diagnosis [1]. However, the energy available per band is significantly reduced because the incoming light is divided into numerous narrow bands. This creates an inherent tradeoff between spatial and spectral resolution. Hyperspectral (HS) images generally provide high spectral resolution (i.e., have many bands) but suffer from low spatial resolution. In contrast, multispectral (MS) images capture high spatial resolution but contain limited spectral information. HS-MS fusion techniques aim to combine HS and MS images of the same scene to produce an enhanced image with improved spatial and spectral resolution [2, 3]. The fused image is particularly useful in applications such as clustering [4] and classification [5]. HS-MS fusion has several practical challenges, including image misregistration, atmospheric variations, and inconsistent illumination [6]. Most studies simplify the problem by assuming perfect registration between HS and MS images and consistent atmospheric and illumination conditions. However, even under these idealized settings, obtaining high-quality fused images necessitates using sophisticated reconstruction algorithms.
@@ 1.1 Literature review
HS-MS fusion techniques can be categorized into five groups: pansharpening [3, 7], Bayesian [8, 9, 10, 11, 12], multiresolution analysis [13], tensor factorization [14, 15, 16, 17, 18, 19], and deep learning [20, 21, 22, 2, 23]. Of these, Bayesian, tensor factorization, and deep learning have garnered attention due to their state-of-the-art performance. In Bayesian methods, the fusion problem is formulated as a maximum a posteriori (MAP) optimization, incorporating appropriate priors on the ground truth. Commonly used priors include Gaussian priors [24], Markov Random Fields [25], transform-domain sparsity priors [12, 19, 8], and low-rank and nonnegativity priors [26]. On the other hand, deep learning methods attempt to learn the functional relationship between the ground truth and the observed HS-MS images using end-to-end deep networks. They can achieve remarkable results, leveraging architectures such as transformers [27, 28, 29, 30, 31], CNNs [32, 33, 34], and diffusion models [35, 36, 37, 38]. Unfortunately, deep learning relies on large training datasets, which can be difficult to obtain when high-resolution ground truth images are unavailable [2]. Training deep models is computationally intensive, and models trained for specific forward models often lack adaptability [2].
Figure 1: An illustration of instability in denoiser-driven fusion on the Pavia dataset. Specifically, we compare the performance of plug-and-play (PnP) fusion using DnCNN [39] and the proposed CasKD denoiser. The plots track the evolution of PSNR and the distance between successive iterates over iterations. The inset images show the fused results at iterations 1, 100, and 1500. Notably, with DnCNN, the PSNR initially improves but then drops sharply with iterations, ultimately leading to a poor reconstruction. This instability is also highlighted in the second plot, showing the divergence of the iterates $\mathbf{X}_{k}$. This phenomenon is, in fact, peculiar to deep denoisers and has been previously observed for other inverse problems [40]. In contrast, the PSNR steadily increases with our denoiser, and the successive differences decay to zero, which aligns with the theoretical guarantees established in this work.
@@ 1.2 Motivation
Recently, state-of-the-art denoisers have been successfully employed for image regularization, achieving state-of-the-art results [41, 42, 43, 44]. Our work focuses on the Plug-and-Play (PnP) method, where an off-the-shelf denoiser is used as a regularizer inside a classical reconstruction algorithm. The denoiser acts as an implicit prior, guiding the iterative process towards a better reconstruction. Compared to end-to-end methods, Plug-and-Play (PnP) allows the decoupling of the forward model and the regularizer. In an end-to-end approach, the forward model is typically utilized during training but not during inference. In contrast, PnP leverages a pretrained denoiser, allowing it to be applied across different forward models. The flip side is that PnP algorithms are strictly not derived from an optimization framework, and the convergence of the iterates is not guaranteed automatically. It is thus not surprising that the convergence of PnP methods has been an active area of research [42, 45, 46, 47, 48, 49]. It is worth noting that the convergence guarantees for deep denoisers often depend on strong technical assumptions that are either impractical or difficult to verify in real-world scenarios.
The choice of the denoiser in PnP can significantly impact the reconstruction quality. Classical denoisers [50, 51, 52, 53, 54, 55, 56] do not require training and offer better interpretability and generalizability. Deep denoisers have recently gained popularity due to their strong denoising capability [39, 57, 58, 53]. In the PnP framework, the denoiser acts more as a regularizer—a smoothing operator—than merely removing “noise”  [59, 46]. Unlike traditional denoising, where the denoiser is applied once to a noisy image, PnP algorithms employ the denoiser iteratively. Consequently, even a relatively weak denoiser can be an effective regularizer, producing a high-quality reconstruction. Rather, the challenge lies in ensuring that the PnP iterations remain stable and do not diverge. As illustrated in Figure 1, powerful deep denoisers such as DnCNN [39] can cause the PnP iterations to become unstable and diverge beyond a point. We have also observed this instability in hyperspectral imaging scenarios, underscoring the need for denoisers that improve image quality and guarantee convergence.
@@ 1.3 Contribution
Recent works have demonstrated that a class of pseudo-linear denoisers, known as kernel denoisers, can be used to develop convergent PnP algorithms [52, 49, 42, 60]. The convergence guarantee stems from the unique mathematical properties of kernel denoisers, which are challenging to establish for (black-box) deep denoisers. While kernel denoisers are not as powerful as DnCNN [39] and DRUNET [61], they have been shown to give good reconstructions for inverse problems such as deblurring and superresolution [62, 52, 42, 48]. Kernel denoisers have also been used for convergent hyperspectral fusion [52]. However, as explained in Section 4, the analysis in these works does not apply to our PnP algorithm. This is because we utilize an enhanced denoiser that lacks specific properties (such as symmetry) required in previous analyses. More precisely, our contribution is as follows.
@@
    Cascaded denoiser. We introduce a hyperspectral denoiser, CasKD, by cascading two existing kernel denoisers [52, 42]. The first denoiser is designed to capture inter-band correlations, leveraging spectral dependencies in hyperspectral images, while the second denoiser exploits in-band correlations. We show that CasKD offers superior denoising and regularization capabilities compared to either denoiser alone, albeit with an increased computational cost. This improvement is significant, as simply cascading two denoisers does not always ensure better performance. We also analyze the mathematical properties of CasKD that are necessary for the convergence analysis.
@@
    Convergence analysis. We integrate the proposed CasKD denoiser with the proximal gradient descent algorithm [63] to develop a fusion algorithm, HyDeFuse. We prove that HyDeFuse is guaranteed to converge linearly to a unique reconstruction for any arbitrary initialization. To the best of our knowledge, there are not many (iterative) fusion algorithms where the iterates are guaranteed to converge at a geometric rate. Importantly, the analysis allows us to identify the correct step size required to ensure convergence.
@@ 1.4 Organization
We review background materials in Section 2, including the forward model and the loss function for HS-MS fusion. The proposed denoiser and the fusion algorithm are presented in Section 3, where we also discuss its mathematical properties. The core part is Section 4, where we analyze the convergence of HyDeFuse. Finally, in Section 5, we validate our theoretical findings and compare the performance of HyDeFuse with existing fusion algorithms.
@@ 2.1 Hyperspectral Imaging
The problem of HS-MS fusion can be formulated as an inverse problem with a linear forward model. This formulation is derived from simplifying assumptions about imaging optics, which inherently include nonlinear distortions, sensor noise, and optical aberrations [64, 65]. The sensor plane in hyperspectral imaging consists of an array of detectors that measure the intensity of incoming light at different wavelengths. In MS imaging, each sensor is sensitive to specific regions of the electromagnetic spectrum. In contrast, HS imaging sensors are designed to capture light across several narrow, contiguous spectral bands. A common approach is to model the sensor plane as a two-dimensional lattice, where the sensors are arranged on a sampling grid [65]. In MS imaging, the sensor distribution is denser for a given ground-truth resolution, leading to a higher spatial sampling rate than in HS imaging. As a result, MS images generally exhibit superior spatial resolution. However, each MS pixel integrates information from multiple spectral bands, resulting in spectral degradation. Although HS images capture finer spectral variations over a broader range, their lower sensor density and narrower spectral bands lead to averaging at the sensor plane, making them appear as blurred versions of the high-resolution ground truth. Moreover, due to the lower sampling rate of HS sensors and the reduced energy received per spectral band, spatial downsampling occurs at the sensor level.
@@ 2.2 Forward Model
The imaging process described above can be modeled using low-pass filtering and downsampling applied to an (unknown) ground truth image with high spatial and spectral resolution [65, 8]. The inverse problem is to estimate this ground truth from the available HS and MS images. They are naturally represented as 3D tensors, with the three dimensions corresponding to spatial rows, columns, and spectral bands. In the literature, these 3D tensors are often reshaped into 2D matrices for easier handling and analysis, where each column represents the vectorized image of a specific spectral band. The standard forward model is given by
$\mathbf{Y}_{h}=\mathbf{A}\mathbf{Z}+\mathbf{\Theta}_{h}\qquad\mbox{and}\qquad\mathbf{Y}_{m}=\mathbf{Z}\mathbf{R}+\mathbf{\Theta}_{m},$ (1)
@@
In the above representation, each column of $\mathbf{Z}$ corresponds to the vectorized form of a spectral band, and the matrices $\mathbf{B},\mathbf{S}$ (resp. $\mathbf{R}$) act on the columns (resp. rows) of $\mathbf{Z}$. They are not explicitly stored as matrices in practice but are applied as operators.
@@ 2.3 Variable Reduction
The simplified forward model (1) reduces the fusion problem to solving a system of linear equations. However, note that the number of equations and variables in (1) are $N_{h}L_{h}+N_{m}L_{m}$ and $N_{m}L_{h}$. Typically, $N_{h}\ll N_{m}$ and $L_{m}\ll L_{h}$, meaning the number of unknowns exceeds the number of measurements, making the computation of $\mathbf{Z}$ a highly ill-posed problem [52]. This is precisely where regularization becomes crucial. Moreover, following prior work [8, 66], we use dimensionality reduction to decrease the number of variables. Specifically, we assume that the bands in $\mathbf{Z}$ are highly correlated and that $\mathbf{Z}$ lies in a subspace of lower dimension. Mathematically, this can be expressed as a decomposition,
@@
where the rows of $\mathbf{E}$ form the basis of the low-dimensional subspace. In other words, the model (1) becomes
$\mathbf{Y}_{h}=\mathbf{A}\mathbf{X}\mathbf{E}+\mathbf{\Theta}_{h}\qquad\mbox{and}\qquad\mathbf{Y}_{m}=\mathbf{X}\mathbf{E}\mathbf{R}+\mathbf{\Theta}_{m},$ (3)
where $\mathbf{X}$ is the unknown (or latent) variable. The reconstructed image is given by $\widehat{\mathbf{Z}}=\widehat{\mathbf{X}}\mathbf{E}$, where $\widehat{\mathbf{X}}$ is obtained by solving (3).
A standard method for estimating $\mathbf{E}$ from the observed image $\mathbf{Y}_{h}$ is using the singular value decomposition [43, 8]. Specifically, we first upsample $\mathbf{Y}_{h}$ to the spatial resolution of $\mathbf{Z}$. The resulting image, say $\mathbf{Y}$, is a surrogate of the ground truth. We perform its singular value decomposition
@@
where $\sigma_{1}\geqslant\cdots\geqslant\sigma_{r}$ are its singular values, and $\boldsymbol{u}_{1},\ldots,\boldsymbol{u}_{r}\in\mathbb{R}^{N_{m}}$ and $\boldsymbol{v}_{1},\ldots,\boldsymbol{v}_{r}\in\mathbb{R}^{L_{h}}$ are the left and right singular vectors, and $r$ is the rank of $\mathbf{Y}$. We choose some $L_{s}\leqslant r$ as the dimension of the targeted subspace and define
$\mathbf{E}=\Big{[}\boldsymbol{v}_{1}\cdots\boldsymbol{v}_{L_{s}}\Big{]}^{\top}\in\mathbb{R}^{L_{s}\times L_{h}}$
as the basis in (3). We will later use the property that $\mathbf{E}\mathbf{E}^{\top}=\mathbf{I}$.
3 Denoiser-Driven Fusion

### v2

# 33footnotetext: K. N. Chaudhury was supported by the ANRF, Government of India, under grant STR/2021/000011

@@ Abstract
Hyperspectral (HS) images provide fine spectral resolution but have limited spatial resolution, whereas multispectral (MS) images capture finer spatial details but have fewer bands. HS-MS fusion aims to integrate HS and MS images to generate a single image with improved spatial and spectral resolution. This is commonly formulated as an inverse problem with a linear forward model. However, reconstructing high-quality images using the forward model alone is challenging, necessitating the use of regularization techniques. In this work, we investigate the paradigm of denoiser-driven regularization, where a powerful off-the-shelf denoiser is used for implicit regularization within an iterative algorithm. This has shown much promise but remains relatively underexplored in hyperspectral imaging. The technical challenge lies in designing hyperspectral denoisers that can guarantee convergence—while strong denoisers can produce high-quality reconstructions, they may also cause instability or divergence. Specifically, we consider a denoiser-driven fusion algorithm, HyDeFuse, which leverages a class of pseudo-linear denoisers for implicit regularization. We demonstrate how the contraction mapping theorem can be applied to establish global linear convergence of HyDeFUse. Finally, we validate our theoretical findings and present fusion results on publicly available datasets to demonstrate the performance of HyDeFuse.
Keywords: hyperspectral fusion, kernel denoiser, plug-and-play method, convergence.
@@ 1 Introduction
Hyperspectral cameras capture images of a scene over a continuous range of spectral bands. Because the incoming light is divided into many narrow bands, the energy in each band decreases, resulting in an inherent trade-off between spatial and spectral resolution. Hyperspectral (HS) images typically offer high spectral resolution (numerous bands) but suffer from low spatial resolution. In contrast, multispectral (MS) images capture high spatial resolution but contain limited spectral information. HS-MS fusion techniques combine HS and MS images of the same scene to generate an image with both high spatial and spectral resolution [1, 2]. Such fused images are useful in applications such as navigation, surveillance, object identification, medical diagnosis, and classification [3, 4, 5].
HS-MS fusion faces several practical challenges, such as misregistrations, atmospheric changes, and inconsistent illumination [6]. Most studies simplify the problem by assuming perfect registration and stable conditions. However, even under these ideal settings, producing high-quality fused images requires advanced reconstruction algorithms.
@@ 1.1 Literature review
HS-MS fusion techniques can be categorized into five groups: pansharpening [2, 7], Bayesian [8, 9, 10, 11, 12], multiresolution analysis [13], tensor factorization [14, 15, 16, 17, 18, 19], and deep learning [20, 21, 22, 1, 23]. Of these, Bayesian, tensor factorization, and deep learning have garnered attention due to their state-of-the-art performance. In Bayesian methods, the fusion problem is formulated as a maximum a posteriori (MAP) optimization, incorporating appropriate priors on the ground truth. Commonly used priors include Gaussian priors [24], Markov Random Fields [25], transform-domain sparsity priors [12, 19, 8], and low-rank and nonnegativity priors [26]. On the other hand, deep learning methods attempt to learn the functional relationship between the ground truth and the observed HS-MS images using end-to-end deep networks. They can achieve remarkable results, leveraging architectures such as transformers [27, 28, 29, 30, 31], CNNs [32, 33, 34], and diffusion models [35, 36, 37, 38]. Unfortunately, deep learning methods require large training datasets, which are often hard to obtain when high-resolution ground truth images are unavailable [1]. Moreover, training deep models is computationally demanding, and models trained for specific forward models often lack adaptability [1].
Figure 1: An illustration of instability in denoiser-driven fusion on the Pavia dataset. The inset images show the fused results at iterations 1, 100, and 1500. Specifically, we compare the performance of Proximal Gradient Descent (PnP-PGD) using DnCNN [39] and the proposed CasKD denoiser. The plots show the evolution of PSNR and the distance between successive iterates $\{\mathbf{X}_{k}\}$. The PSNR evolution serves as an indicator of stability: ideally, it should improve gradually rather than rise initially and then degrade after many iterations. Notably, with DnCNN, the PSNR initially improves but then starts dropping after a certain point, ultimately leading to a poor reconstruction. The instability is also highlighted in the second plot, showing the divergence of the iterates. This phenomenon is common to deep denoisers and has been previously observed for other inverse problems [40]. In contrast, the PSNR for our denoiser increases steadily. Moreover, the successive differences decay to zero, which aligns with the theoretical guarantees established in this work.
@@ 1.2 Motivation
HS–MS fusion is closely related to the classical problem of image superresolution in computational imaging. Recently, denoisers have been successfully incorporated into such imaging tasks, yielding state-of-the-art results [41, 42, 43, 44]. Our work focuses on the Plug-and-Play (PnP) method, where an off-the-shelf denoiser is used as a implicit regularizer inside a classical reconstruction algorithm. The denoiser acts as an prior, guiding the iterative process towards a better reconstruction. In end-to-end approaches, the forward model is incorporated during training, which means that any change to the forward model necessitates retraining. In contrast, PnP allows the decoupling of the forward model and the regularizer, so that the same pretrained denoiser can be used across different models. The choice of the denoiser in PnP can significantly impact the reconstruction quality. Classical denoisers [45, 46, 47, 48, 49, 50, 51] do not require training and offer better interpretability and generalizability. Deep denoisers have recently gained popularity due to their strong denoising capability [39, 52, 53, 48]. In the PnP framework, the denoiser functions as an implicit regularizer or smoothing operator rather than serving solely as a noise-removal tool [54, 55].
The flip side is that PnP algorithms are not strictly derived from an optimization formulation, and hence the convergence of their iterates cannot be guaranteed automatically. It is thus not surprising that the convergence of PnP methods has been an active area of research [56, 57, 58, 42, 59, 55, 60, 61, 62]. Unlike traditional denoising, where the denoiser is applied once to a noisy image, PnP algorithms employ the denoiser iteratively. Consequently, even a relatively weak denoiser can be an effective regularizer, producing a high-quality reconstruction. Instead, the challenge lies in ensuring that the PnP iterations remain stable and do not diverge from their intended path. By “stable,” we refer to the consistency of PSNR, which should improve steadily rather than rise initially and then degrade after many iterations. As shown in Figure 1, powerful deep denoisers such as DnCNN [39] can exhibit this type of instability, with the iterates diverging after a certain point. We have also observed similar behavior in hyperspectral imaging, highlighting the need for denoisers that ensure convergence or, at the very least, maintain PSNR stability. In this context, it is worth noting that existing convergence guarantees for deep denoisers typically rely on strong technical assumptions, which are often impractical or difficult to verify in real-world settings.
@@ 1.3 Contribution
Recent works have demonstrated that a class of pseudo-linear denoisers, known as kernel denoisers, can be used to develop convergent PnP algorithms [47, 62, 42, 63]. The convergence guarantee stems from the unique mathematical properties of kernel denoisers, which are challenging to establish for (black-box) deep denoisers. While kernel denoisers are not as powerful as DnCNN [39] and DRUNET [64], they have been shown to yield good reconstructions for inverse problems such as deblurring and superresolution [65, 47, 42, 61]. Kernel denoisers have also been used for convergent hyperspectral fusion [47]. However, as explained in Section 4, the analysis in these works does not apply to our PnP algorithm. This is because we utilize an enhanced denoiser that lacks specific properties (such as symmetry) required in previous analyses. Our contributions are as follows:
@@
    Cascaded denoiser. We introduce a hyperspectral denoiser, CasKD, by cascading two existing kernel denoisers [47, 42]. The first denoiser is designed to capture inter-band correlations, leveraging spectral dependencies in hyperspectral images, while the second denoiser exploits the in-band correlations. We demonstrate that CasKD provides superior denoising and fusion capabilities compared to either denoiser alone, although at an increased computational cost. This improvement is significant, as simply cascading two denoisers does not always ensure better performance. We also analyze the mathematical properties of CasKD that are necessary for the convergence analysis.
@@
    Convergence analysis. We integrate the proposed CasKD denoiser with the proximal gradient descent algorithm [66] to develop a fusion algorithm called HyDeFuse. We prove that HyDeFuse is guaranteed to converge linearly to a unique reconstruction for any arbitrary initialization. To the best of our knowledge, there are not many (iterative) fusion algorithms where the iterates are guaranteed to converge at a geometric rate. Importantly, the analysis enables us to determine the optimal step size required for convergence.
@@ 1.4 Organization
We review background materials in Section 2, including the forward model and the loss function for HS-MS fusion. The proposed denoiser and the fusion algorithm are presented in Section 3, where we also discuss their mathematical properties. The core part is Section 4, where we analyze the convergence of HyDeFuse. Finally, in Section 5, we validate our theoretical findings and compare the performance of HyDeFuse with existing fusion algorithms. Proofs of technical results are provided in Section 7.
@@ 2.1 Hyperspectral Imaging
The problem of hyperspectral–multispectral (HS–MS) fusion can be posed as an inverse problem governed by a linear forward model. This formulation arises from simplified assumptions about the imaging optics, which in reality involve nonlinear distortions, sensor noise, and optical aberrations [67, 68]. In hyperspectral imaging, the sensor plane comprises an array of detectors that record the intensity of incoming light at various wavelengths. It is often modeled as a two-dimensional lattice, where sensors are positioned on a uniform sampling grid [68].
In MS imaging, each sensor is responsive to a specific region of the electromagnetic spectrum. For a given ground-truth resolution, MS sensors are more densely distributed, resulting in a higher spatial sampling rate than in HS imaging. Consequently, MS images typically offer superior spatial resolution. However, since each MS pixel integrates information from multiple spectral bands, this comes at the cost of spectral degradation.
In contrast, HS sensors capture light over several narrow and contiguous spectral bands. Although this enables fine spectral discrimination across a wide range, the lower sensor density and narrower bands lead to spatial averaging at the detector plane, making HS images appear as blurred representations of the high-resolution scene. Additionally, the reduced sampling rate and limited energy per band result in spatial downsampling at the sensor level.
@@ 2.2 Forward Model
The above imaging process can be modeled as a combination of low-pass filtering and downsampling applied to an unknown high-resolution image with fine spatial and spectral details [68, 8]. The corresponding inverse problem aims to recover this high-resolution image from the observed HS and MS measurements.
The unknown image is naturally represented as a 3-D tensor, whose dimensions correspond to spatial rows, spatial columns, and spectral bands. For computational convenience, this 3-D tensor is often reshaped into a matrix, where each column corresponds to the vectorized image of a specific spectral band. Under this convention, the standard forward model can be expressed as
$\mathbf{Y}_{h}=\mathbf{A}\mathbf{Z}+\mathbf{\Theta}_{h},\qquad\mathbf{Y}_{m}=\mathbf{Z}\mathbf{R}+\mathbf{\Theta}_{m},$ (1)
@@
In the above representation, each column of $\mathbf{Z}$ corresponds to the vectorized form of a spectral band, and the matrices $\mathbf{B},\mathbf{S}$ (resp. $\mathbf{R}$) act on the columns (resp. rows) of $\mathbf{Z}$. In practice, the matrices $\mathbf{B}$, $\mathbf{S}$, and $\mathbf{R}$ are not stored explicitly but are implemented as linear operators.
@@ 2.3 Variable Reduction
The simplified forward model in (1) reduces the fusion task to solving a system of linear equations. The total number of equations and unknowns in (1) are $N_{h}L_{h}+N_{m}L_{m}$ and $N_{m}L_{h}$, respectively. Typically, $N_{h}\ll N_{m}$ and $L_{m}\ll L_{h}$, implying that the number of unknowns exceeds the number of measurements. As a result, estimating $\mathbf{Z}$ becomes a highly ill-posed problem [47], thereby necessitating regularization.
To further reduce the number of variables, we follow prior work [8, 69] and employ dimensionality reduction. The key assumption is that the spectral bands in $\mathbf{Z}$ are highly correlated, and hence $\mathbf{Z}$ lies in a low-dimensional subspace. Mathematically, this can be expressed as
@@
where the rows of $\mathbf{E}$ form a basis for the low-dimensional subspace. Substituting (2) into (1) yields
$\mathbf{Y}_{h}=\mathbf{A}\mathbf{X}\mathbf{E}+\mathbf{\Theta}_{h},\qquad\mathbf{Y}_{m}=\mathbf{X}\mathbf{E}\mathbf{R}+\mathbf{\Theta}_{m},$ (3)
where $\mathbf{X}$ is the variable to be estimated. The final reconstruction is then given by $\widehat{\mathbf{Z}}=\widehat{\mathbf{X}}\mathbf{E}$, where $\widehat{\mathbf{X}}$ denotes the solution of (3).
A standard approach for estimating $\mathbf{E}$ from the observed hyperspectral image $\mathbf{Y}_{h}$ is to use singular value decomposition (SVD) [43, 8]. Specifically, $\mathbf{Y}_{h}$ is first upsampled to match the spatial resolution of $\mathbf{Z}$, and the resulting image, denoted by $\mathbf{Y}$, serves as a surrogate for $\mathbf{Z}$. We then compute its SVD as
@@
where $r$ is the rank of $\mathbf{Y}$, $\sigma_{1}\geqslant\cdots\geqslant\sigma_{r}$ are the singular values, and $\boldsymbol{u}_{1},\ldots,\boldsymbol{u}_{r}\in\mathbb{R}^{N_{m}}$ and $\boldsymbol{v}_{1},\ldots,\boldsymbol{v}_{r}\in\mathbb{R}^{L_{h}}$ are the left and right singular vectors, respectively.
Following the approach in [8], we select a target subspace dimension $L_{s}\leqslant r$ and define
$\mathbf{E}=\big[\boldsymbol{v}_{1}\ \cdots\ \boldsymbol{v}_{L_{s}}\big]^{\top}\in\mathbb{R}^{L_{s}\times L_{h}},$
which serves as the basis of the low-dimensional subspace in (3). The dimension $L_{s}$ is chosen such that most of the signal energy is preserved after projection onto the subspace. If $L_{s}$ is too small, important spectral information may be lost. In practice, we found that the performance is not highly sensitive to the exact value of $L_{s}$. A smaller $L_{s}$ also reduces computational cost per iteration, since fewer spectral bands are processed. Hence, it is preferable to choose the smallest $L_{s}$ that still retains most of the signal energy. The specific choice of $L_{s}$ is discussed in Section 5.1.